OpenStudy (mathmath333):

973 – 433 – 543 is always divisible by: a.)43 b.)54 c.)97 d.)all of the above

OpenStudy (fibonaccichick666):

So can you compute 973 – 433 – 543?

OpenStudy (mathmath333):

its -3

OpenStudy (fibonaccichick666):

alright, so I see your dilemma, can you check and make sure you copied the question correctly?

OpenStudy (fibonaccichick666):

if not, we shall have to approach this differently

OpenStudy (mathmath333):

yes its 100% correct lol

OpenStudy (fibonaccichick666):

okie dokey, so do you remember the distributive property?

OpenStudy (mathmath333):

a(b+c)=ab+ac ?

OpenStudy (fibonaccichick666):

yea, so we can extend that to 3 numbers like so a(b+c+d)=ab+ac+ad

OpenStudy (fibonaccichick666):

so, if we figure out which numbers divide all 3 terms evenly, we can determine our answer

OpenStudy (fibonaccichick666):

make sense?

OpenStudy (fibonaccichick666):

that is accurate, but it is not the distributive property

OpenStudy (fibonaccichick666):

distributive property is like (12+18+21)=3(4+6+7)

OpenStudy (mathmath333):

oh wait it should be \(\large 973 – 433 – 543=970-430-540-3\)

OpenStudy (fibonaccichick666):

no

OpenStudy (fibonaccichick666):

it'd be like \[3(324.333+144.333+181)=973-433-543\]

OpenStudy (fibonaccichick666):

if you used 3.

OpenStudy (fibonaccichick666):

Do you see that I divided?

OpenStudy (mathmath333):

ok but why r u taking 3 commn

OpenStudy (fibonaccichick666):

commn? I don't know that word. And I used 3 because that is the number you were attempting to use above

OpenStudy (fibonaccichick666):

I wanted to show you it does not work

OpenStudy (mathmath333):

->common

OpenStudy (fibonaccichick666):

ah sorry, I'm tired lol You're my last question for the night

OpenStudy (fibonaccichick666):

but anyways, do you understand that when you factor something out of parentheses you will be dividing each term in the parentheses by that number?

OpenStudy (mathmath333):

i do

OpenStudy (fibonaccichick666):

ok, so how do you think we can arrive at our answer given this information?

OpenStudy (mathmath333):

i d k that's y i posted

OpenStudy (fibonaccichick666):

Ya know what, that way doesn't even work, none of these solutions work O.O or I am just that tired. I'm gonna hand off to @ganeshie8 or @ParthKohli as I do not think that the solution exists the way the question is phrased.

OpenStudy (fibonaccichick666):

are you sure that last answer isn't None of the above?

Parth (parthkohli):

my first thought was maybe it would be asking about divisibility of individual numbers, but that wouldn't work too.

OpenStudy (mathmath333):

let me post a screenshot

OpenStudy (mathmath333):

http://prntscr.com/62h05w

ganeshie8 (ganeshie8):

-3 is divisibly by +-1 and +-3, they are not there in the options so the question is wrong

OpenStudy (fibonaccichick666):

yea, the question has a type-o

OpenStudy (fibonaccichick666):

I thought I was just having a dumb day

Parth (parthkohli):

maybe look at order of operations too... 973 - (433 - 543) none of the options work here either

OpenStudy (mathmath333):

still have a hope for a miracle

Parth (parthkohli):

conclusion: don't give a damn about that website

OpenStudy (fibonaccichick666):

It should be 430 or 540 then the question would make sense